OpenStudy (anonymous):

Find the volume of the region bounded in back by the plane x=0 , on the front and sides by the parabolic cylinder x=1−y2 , on top by the paraboloid z=x2+y2 , and on the bottom by the xy-plane. Find the volume of the region bounded in back by the plane x=0 , on the front and sides by the parabolic cylinder x=1−y2 , on top by the paraboloid z=x2+y2 , and on the bottom by the xy-plane. @Mathematics

OpenStudy (turingtest):

Bounded in back by \(x=0\) and in front by \(1-y^2\)\[0\le x\le1-y^2\]this suggeests that\[0\le y\le1\]since it is bounded below by the xy-plane, \(z\ge0\) so we have\[0\le z\le x^2+y^2\]I think from here we want to convert this to cylindrical coordinates